OpenStudy (anonymous):

Consider the function f(x) = 6sin(x2) on the interval 0 ≤ x ≤ 3. (a) Find the exact value in the given interval where an antiderivative, F, reaches its maximum. x = sqrt(pi) (b) If F(1) = 5, estimate the maximum value attained by F. (Round your answer to three decimal places.) … I just need help with part b; I don'y know how to begin

OpenStudy (anonymous):

its f(x)=6sin(x^2) not "x2"… sorry :/

OpenStudy (freckles):

so the first question sounds like you want to find \[\frac{d}{dx}\int\limits_{0}^{x}f(t) dt=0\] this will help you find the critical numbers of F

OpenStudy (freckles):

find when x satisfies the equation:*

OpenStudy (freckles):

basically solve 6sin(x^2)=0 for x

OpenStudy (anonymous):

i did i got sort (pi) for a

OpenStudy (anonymous):

sqrt*

OpenStudy (freckles):

I guess you already did a I got 2 zeros to 6sin(x^2)=0 on [0,3] then you can plug them into the derivative of 6sin(x^2) to see if it is a max or not

OpenStudy (anonymous):

so it'd be -6cos(x^2)*2x = -12cos(x^2)? I just plug the two 0's into that and see what the max value is right..?

OpenStudy (freckles):

where you get the negative at?

OpenStudy (freckles):

\[12xcos(x^2)\]

OpenStudy (anonymous):

oh my bad its just cos, i got confused with the -cos

OpenStudy (anonymous):

i was thinking about antiderivatives

OpenStudy (freckles):

yep one of the zeros gives a min and one of them gives a max

OpenStudy (freckles):

and you chose the right zero that will give us a max gj

OpenStudy (anonymous):

for 0 i got 0 and for 3 i got -32.80068943

OpenStudy (anonymous):

and for sqrt(pi) i got -21.26944621

OpenStudy (freckles):

\[6 \sin(x^2)=0 \\ \sin(x^2)=0 \\ x^2=n \pi\] where n is integer greater than 0 we want x to be between 0 and 3 so we want x^2 to be between 0 and 9 so n could be 1 or 2 xo we x=sqrt(pi) or x=sqrt(2pi)

OpenStudy (freckles):

but 12sqrt(pi)cos(pi) is negative so at x=sqrt(pi) we have a max and 12sqrt(2pi)cos(2pi) is positive so at x=sqrt(2pi) we have a min

OpenStudy (freckles):

so it want's us to use F(1)=5 to estimate F(sqrt(pi))

OpenStudy (anonymous):

thats where i don't know what to do:/

OpenStudy (freckles):

\[F(x)=\int\limits_{0}^{x}6 \sin(t^2) dt \] we have \[F(1)=\int\limits_{0}^{1}6 \sin(t^2) dt =5 \\ F(\sqrt{\pi})=\int\limits_{0}^{\sqrt{\pi}}6 \sin(t^2) dt=n\] n being the value we want to obtain...

OpenStudy (anonymous):

so finding the area under the curve of f(x) using the interval given will give me the value of n..?

OpenStudy (freckles):

we have have (1,5) is on F we can find the slope at (1,5) of F

OpenStudy (freckles):

\[y-y_1=m(x-x_1) \\ y=m(x-x_1)+y_1 \] my idea is to maybe a build a tangent line to F at (1,5)

OpenStudy (anonymous):

hmm okay

OpenStudy (freckles):

you already found the derivative of F use that to find the slope at x=1

OpenStudy (anonymous):

6.48362767

OpenStudy (freckles):

is that the slope?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

I got 6sin(1) is approximately 5.0488

OpenStudy (anonymous):

oh because i plugged in the value into the derivative of the function, not the original function

OpenStudy (freckles):

\[F(x) \approx 5.0488(x-1)+5 \\ \text{ this is usually a pretty good approximation for values near x=1 } \\ \]

OpenStudy (freckles):

well F'=f right?

OpenStudy (freckles):

the derivative of big F is small f

OpenStudy (anonymous):

true!!!

OpenStudy (freckles):

I'm just worried about how good that approximation will be since sqrt(pi) is a little further than 1 than I would like

OpenStudy (freckles):

but I can't see how to get a better approximation right now

OpenStudy (anonymous):

i see what you're saying, and exactly what i was thinking...

OpenStudy (freckles):

however if we do use that "approximation" we will get 8.89997

OpenStudy (freckles):

Have you learned any other approximation techniques?

OpenStudy (anonymous):

i am afraid not

OpenStudy (freckles):

sounds like the linear approximation is the way to go then

OpenStudy (anonymous):

okay i will try that!

OpenStudy (freckles):

is this an online assignment?

OpenStudy (freckles):

well i think it said round to 3 decimal places

OpenStudy (anonymous):

yes it was wrong :/ its a tough question

OpenStudy (freckles):

https://answers.yahoo.com/question/index?qid=20101120210256AAl0I6z this person did this...

OpenStudy (freckles):

I don't know what technique that is

OpenStudy (freckles):

they had F(1)=7

OpenStudy (anonymous):

one of my classmates had that and they inputed that answer and it was wrong as well

OpenStudy (freckles):

did you enter in 9?

OpenStudy (anonymous):

I was going to try that next haha

OpenStudy (freckles):

because 8.89997 to 3 decimal poitns shuold be 9.000 or just 9

OpenStudy (anonymous):

9 was wrong as well

OpenStudy (freckles):

you get infinite amount of answers?

OpenStudy (anonymous):

hmm i think I'm going to move onto another question… will you be willing to help me with another question..? Im a bit lost in this chapter because i wasn't able to attend the lecture :/

OpenStudy (freckles):

lol try this 1.169+5

OpenStudy (freckles):

6.169

OpenStudy (anonymous):

haha okay i'll give it a shot lol

OpenStudy (anonymous):

lol it was wrong too

OpenStudy (anonymous):

ill post the next question

OpenStudy (freckles):

I'm sorry I put in the wrong integral into wolfram I put 2sinx^2

OpenStudy (freckles):

\[\int\limits\limits_{1}^{\ \sqrt{\pi}}6 \sin(x^2)+F(1) \\ 3.507+5=8.507\]

OpenStudy (freckles):

not too far off from our approximation